1.It is claimed that the average weight of babies at birth is 3.4 kg. The average weight of a random sample of 30 newly born babies was determined.It was found out that the average weight was 3.1 kg. Is there a reason to believe that the average weight of babies at birth is not 3.4 kg? Assume that the population standard deviation is 1.1kg. Use 0.05 level of significance.



1
Expert's answer
2022-03-01T04:38:40-0500

The following null and alternative hypotheses need to be tested:

"H_0:\mu=3.4"

"H_1:\mu\not=3.4"

This corresponds to a two-tailed test, for which a t-test for one mean, with unknown population standard deviation, using the sample standard deviation, will be used.

Based on the information provided, the significance level is "\alpha = 0.05," "df=n-1=29" degrees of freedom, and the critical value for a two-tailed test is "t_c = 2.04523."

The rejection region for this two-tailed test is "R = \{t: |t| > 2.04523\}."

The t-statistic is computed as follows:


"t=\dfrac{\bar{X}-\mu}{s/\sqrt{n}}=\dfrac{3.1-3.4}{1.1/\sqrt{30}}=-1.49379"

Since it is observed that "|t| = 1.49379\le 2.04523=t_c," it is then concluded that the null hypothesis is not rejected.

Using the P-value approach:

The p-value for two-tailed, "df=29" degrees of freedom, "t=-1.49379," is "p=0.146032," and since "p=0.146032>0.05=\alpha," it is concluded that the null hypothesis is not rejected.

Therefore, there is not enough evidence to claim that the population mean "\mu" is different than "3.4," at the "\alpha = 0.05" significance level.